OpenStudy (anonymous):

What is the mean of the following set of data? 12, 16, 10, 30 12 17 10 30

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

add them all up then divide them by how many numbers they are. so __/4=?

OpenStudy (anonymous):

so what is 30+10+12+16

OpenStudy (anonymous):

17

OpenStudy (anonymous):

so 30+10+12+16=68

OpenStudy (anonymous):

68/4= what

OpenStudy (anonymous):

68/4=17

OpenStudy (anonymous):

68

OpenStudy (anonymous):

17 you got it

OpenStudy (anonymous):

if you ever need help with median, mean, and the other m word just @ me